OpenStudy (anonymous):

Complete the inverse of F(x) below. Do not include "F-1(y) =" in your answer F(x) x/3 - 1/3 f^-1(y)=???

OpenStudy (binary3i):

\[y=x/3 -1/3\] \[3(y+1/3)=x\] now put y inplace of x and x in place of y.

OpenStudy (binary3i):

it becomes y=3x+1

OpenStudy (anonymous):

is that the final answer? or do more with it?

OpenStudy (campbell_st):

swap x and y \[x = \frac{y}{3} - \frac {1}{3}\] now make y the subject add 1/3 to both sides \[x + \frac{1}{3} = \frac{y}{3}\] multiply both sides by 3 \[y = 3x + 1 \] then \[f^{-1}(x) = 3x + 1\]